OpenStudy (anonymous):

Ten less than four times a certain number is 18. Determine the number.

OpenStudy (anonymous):

with soultion and checking

OpenStudy (anonymous):

let x= certain number 4x-10=18 Because it is said that 10 less than four times the certain number (x) is 18. Then transpose and solve the operations 4x=18+10 Divide both sides by 4 4x=28 x=7 If we check, 4(7)-10= 18 18=18 :D